OpenStudy (anonymous):

2a^2 - 4a + 2 / 3a^2 - 3

jimthompson5910 (jim_thompson5910):

The problem is \[\Large \frac{2a^2 - 4a + 2}{3a^2-3}\] right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

alright, thanks

jimthompson5910 (jim_thompson5910):

We can factor 2a^2 - 4a + 2 to get 2a^2 - 4a + 2 2(a^2 - 2a + 1) 2(a-1)(a-1)

jimthompson5910 (jim_thompson5910):

Then we can factor 3a^2-3 to get 3a^2-3 3(a^2-1) 3(a-1)(a+1)

jimthompson5910 (jim_thompson5910):

So \[\Large \frac{2a^2 - 4a + 2}{3a^2-3}\] turns into \[\Large \frac{2(a-1)(a-1)}{3(a-1)(a+1)}\]

jimthompson5910 (jim_thompson5910):

Do you see where to go from here?

OpenStudy (anonymous):

yes thanks

jimthompson5910 (jim_thompson5910):

you're welcome